# Robust Few-shot Transfer Learning for Knowledge Base Question Answering with Unanswerable Questions

### v1

@@ 5 Conclusions
For robust, low-resource KBQA, we have proposed the practical task of few-shot transfer learning with unanswerability. We have proposed FUn-FuSIC that adapts the existing SoTA model that assumes answerability by (i) candidate logical form generation by iterative LLM prompting with feedback from diverse error checks – including a novel back-translation based equivalence check, and (ii) assessing confidence in the candidates by adapting self consistency to detect answerability and different categories of unanswerability. Using newly created datasets, we show that FUn-FuSIC outperforms adaptations of SoTA models for this setting, and also for answerable KBQA. Our error analysis suggests that performing well across all categories of unanswerability in the few-shot transfer setting is still a challenge and should be a focus of further research.
@@ 7 Risks
At the highest level, our work reduces risk compared to existing KBQA systems, which when improperly adapted in a low-resource setting, incorrectly answer unanswerable questions, without admitting lack of knowledge. However, can incorrectly inferring unaswerability, citing lack of knowledge when the knowledge is in fact available, generate a new category of risk? While we cannot imagine such a risk at the present time, this may require more thought. In any case, KBQA models for unanswerability should strive to minimize this type of error, along with the other types.
@@ A.1.1 KBQA:
A.1.2 Few Shot Transfer Learning for KBQA:
@@
A.1.3 KBQA with unanswerability:
@@
Unanswerable questions in KBQA can be categorized into (a) Schema Level Unanswerability : the question does not have a corresponding logical form that is valid for the KB , (b) Data level unanswerability: it has a valid logical form $l$ for the KB, but which on executing returns an empty answer. Schema level unanswerable questions can further be categorized into (1) Missing Class: The class/type required to construct the logical form is not defined for the KB, (2) Missing Relation: The relation required to construct the logical form is not defined for the KB, (3) Missing Entity: The topic entity specified in the question is missing from the KB. Data level unanswerable questions can be categorized into (1) Missing entity: all classes and relations required to construct the logical form are present in the KB , but there exists no path from the topic entity node to the answer node in the KB due to missing intermediated entities (2) Missing Fact: all classes, relations and entities required to answer the question are present in the KB. However, the (subject, relation, object) path is not connected in the KB.
A.1.4 Few-shot Transfer Learning for KBQA with Unanswerability:
@@
This setting is a generalisation of few shot transfer learning for KBQA, wherein we no longer assume that the test question is answerable with respect to the Knowledge Base $G^{t}$. The training set $D^{s}$ contains thousands of answerable questions along with their corresponding logical form. The target few shots $D^{t}$ contain both answerable as well as unanswerable questions, along with their corresponding logical forms, if they exist.
A.2 Algorithm
Figure 2: Feedback Algorithm for Robustness in KBQA
1:Given: KB-Engine, KB-Ontology, qn $\triangleright$ qn is natural language query
@@ A.2 Algorithm
3:$L\leftarrow[]$
4:for $i=0$ to $N$ do
5:     $lf\leftarrow LLM(prompts)$
6:     $syntax\_error\leftarrow KB-Engine(lf)$
7:     if $syntax\_error\neq\text{null}$ then
8:         $syntax\_prompt\leftarrow SynPrompt(lf,syntax\_error)$
9:         $prompts.append(syntax\_prompt)$
10:         continue
11:     end if
12:     $kb\_consistency\_error\leftarrow CheckTypes(lf,KB-Ontology)$
13:     if $kb\_consistency\_error\neq\text{null}$ then
14:         $kb\_consistency\_prompt\leftarrow LFPrompt(lf,kb\_consistency\_error)$
15:         $prompts.append(kb\_consistency\_prompt)$
16:         continue
17:     end if
18:     $qlf\_error\leftarrow CheckMeaning(lf,qn)$
19:     $ans\_consistency\_error\leftarrow CheckAnswer(KB-Engine(lf),lf.nodes)$
20:     if $qlf\_error\neq\text{null}$ then
21:         $qlf\_prompt\leftarrow NLPrompt(lf,qlf\_error)$
22:         $prompts.append(qlf\_prompt)$
23:         if $ans\_consistency\_error$ = null then
24:              $L.append(lf)$
25:         end if
26:         continue
27:     end if
28:     if $ans\_consistency\_error\neq\text{null}$ then
29:         $ans\_consistency\_prompt\leftarrow EGF(lf,egf\_error)$
30:         $prompts.append(ans\_consistency\_prompt)$
[…]
32:              $L.append(lf)$
33:         end if
34:         continue
35:     end if
36:     return $lf$triangleright all checks passed
37:end for
38$\triangleright$ despite N-tries, no lf passes
39:$L_{ne}\leftarrow[l\text{ for }l\text{ in }L\text{ if }KB-Engine(l)\neq\text{null}]$
40:$L_{e}\leftarrow[l\text{ for }l\text{ in }L\text{ if }KB-Engine(l)=\text{null}]$
41:$poss\_answers\leftarrow[KB-Engine(lf)\text{ for }lf\text{ in }L_{ne}]$
42:if $scUn(poss\_answers)\neq\text{null}$ then $\triangleright$ No Answer-aware-self-consistency
43:     return $SelectBestLF(qn,[lf\text{ for }lf\text{ in }L_{ne}\text{ if }KB-Engine(lf)==scUn(poss\_answers)])$
44:else if $is\_unanswerable$ then
45:     if $L_{e}\neq[]$ then
46:         return $SelectBestLF(qn,L_{e})$
47:     end if
48:     return NK
49:else
50:     return $SelectBestLF(qn,[lf\text{ for }lf\text{ in }L_{ne}\text{ if }KB-Engine(lf)==sc(poss\_answers)])$triangleright self-consistency, with tie breaking
51:end if
Algorithm 1 Self-Consistency Algorithm
1:Given: $possible\_answers$
2:Filter out null values from possible_answers:
3:$possible\_answers\leftarrow[p\text{ for }p\text{ in }possible\_answers\text{ if }p\neq\text{null}]$
4:if $possible\_answers=[]$ then
5:     return null
6:end if
7:$MajVote\leftarrow MaxFreqItem(possible\_answers)$
8:return $MajVote$
Algorithm 2 Self-Consistency with Unanswerability Algorithm
1:Given: $possible\_answers$
2:$MajVote\leftarrow self-consistency(possible\_answers)$
3:if $MajVote=\text{null}$ then
4:     return null
5:end if
6:$count\leftarrow Cnt(MajVote,possible\_answers)$
7:if $count>\frac{\text{len}(possible\_answers)}{2}$ then
8:     return $MajVote$
9:else
10:     return null
11:end if
A.3 Working Example of The Algorithm
@@
Appendix B Ablations
We perform ablations of our best performing model - FUn-FuSIC by removing different types of feedback from the model. We report performance on 100 questions(50 answerable + 50 unanswerable) sampled from the WebQSP $\rightarrow$ GraphQAbility dataset. Refer to 5
Ablation WebQSP $\rightarrow$ GraphQAbility
Overall Ans Unans
F1(L) SM F1 SM F1(L) F1(R) SM
FUn-FuSIC 74.5 54.0 57.0 46.0 92.0 90.0 62.0
sc 64.5 44 63.0 52.0 66.0 64.0 36.0
w/o syntax 74.0 54.0 57.0 46.0 92.0 90.0 62.0
w/o kb-inc 71.8 43.0 51.7 42.0 92.0 88.0 44.0
w/o qf-lf 51.7 24.0 47.7 38.0 55.8 48.0 10.0
w/o ans-inc 73.5 53.0 55.0 44.0 92.0 90.0 62.0
Table 5: Detailed ablation Performance of FUn-FuSIC on WebQSP $\rightarrow$ GraphQAbility. sc denotes replacing self consistency for unaswerability with standard self consistency, while other rows indicate removing check for syntax error (syntax), KB inconsistency (kb-inc), question logical form disagreement (q-lf) and answer incompatibility (ans-inc).
Appendix C EM-s: EM for sparql
Appendix D Performance across different categories of unanswerability
Appendix E Cost Analysis
Hence, the two models are comparable in terms of cost.
Appendix F FuSIC-KBQA: High Level Overview
Our proposed approach FUn-FuSIC builds upon the the base architecture of FuSIC-KBQA. FuSIC-KBQA has a three step pipeline: (a) Supervised Retrieval: a supervised retriever, trained on the source domain and optionally fine-tuned on the target domain is used to obtain the top-100 classes, relations and paths that are relevant to the question asked, (b) LLM ReRanking: an LLM is used to select the top-10 classes, top-10 relations and top-5 paths among the top-100 elements as ranked by the supervised retriever, (c) LLM Generation: We provide the top-10 classes, top-10 relations and top-5 paths along with few-shot exemplars to generate the sparql query.
Appendix G Prompts
G.1 PUn prompt
Header Prompt Translate the following question to sparql for Freebase based on the candidate sparql, candidate entities, candidate relations and candidate entity types which are separated by "|" respectively. Please do not include any other relations, entities and entity types. Your final sparql can have three scenarios: 1. When you need to just pick from candidate sparql. 2. When you need to extend one of candidate sparql using the candidate relations and entity types. 3. When you will generate a new sparql only using the candidate entities, relations and entity types.For entity type check please use this relation "type.object.type".Do not use entity names in the query. Use specified mids. If it is impossible to construct a query using the provided candidate relations or types, return "NK"Make sure that the original question can be regenerated only using the identified entity types, specific entities and relations.
G.2 FUN prompt
syntax error feedback Correct the syntax of the following sparql query. Return ONLY the corrected sparql query without any explanation sparql: SELECT ?x AND ?y ... Virtuoso error: word AND not defined
Target KB Disagreement feedback The generated sparql has a semantic issue warning: The types of relations don’t match for variable ?x in the query. The assigned relation types by [’computer.computer_emulator.computer’, ’type.object.type computer.computer_peripheral’] are [’computer.computer’, ’computer.computer_peripheral’]. These types are mutually incompatible... Please generate again a different executable sparql using the same context and constraints. DO NOT APOLOGIZE - just return the best you can try.
Question Logical form disagreement feedback The question that you answer is NOT same as what you’ve been asked for! You have answered the question "Which opera productions has Gino Marinuzzi conducted?" but you were asked to answer "what is the name of the premiere opera production conducted by gino marinuzzi?". Please generate again a different executable sparql using the relations, classes and entities provided earlier. DO NOT APOLOGIZE - just return the best you can try.
empty answer feedback The generated sparql gives an empty answer when executed on freebase KG, Please generate again a different executable sparql using the same context and constraints.
intermediate node feedback The generated sparql returns an intermediate type node when executed on the freebase KG. Maybe the answer node is an adjacent node to what we currently query for. Please generate again a different executable sparql using the same context and constraints.
answer node in question feedback The logical form upon execution returns International System of Units, which is not answering the question. Please reconstruct the query using same context and constraints.
G.3 Question Logical Form Agreement Prompt
Naturalization of variable names change the sparql query to have variable names representative of what objects they refer to. transform the variable names in this query. Do NOT change the prefix headers and relation names
@@ G.3 Question Logical Form Agreement Prompt
Some examples of few shots are provided below-
question logical form agreement check Question we answer: Who are the cricket players who have made exactly 31 stumps in one day internationals? Question originally asked: name the cricket player who has 31 odi stumps. explanation: The question we answer returns cricket players. The question originally asked also returns cricket players. The question we answer finds cricket players who have made exactly 31 stumps in one day internationals. The question originally asked also finds cricket players who have made 31 stumps in one day internationals. Both questions involve no mathematical or logical operators. Hence, they are same. Question we answer: Which game engines are successors to the Unreal Engine? Question originally asked: which video game engine’s successor engine is unreal engine? explanation: The question we answer returns game engines. The question originally asked also returns game engines. The question we answer finds successors to the Unreal Engine. The question originally asked finds the predecessor of the Unreal Engine. The reasoning steps followed by the two questions are different. Hence, they are different. [total 6 exemplars] Question we answer: Which cars drive at a speed of 80? Question originally asked: name the car with driving speed atleast 80? explanation:
G.4 Logical Form Selection Prompt
We store the back-translated natural language questions along with each predicted logical form. Here, we prompt the LLM to select the closest back-translated natural language question to the orginal question.
Best Logical Form Selection Prompt orig_nl_qn = which surf films has sarah finn served as the casting director? 1. pred_nl: Which surfing films has Sarah Finn directed the casting for? 2. pred_nl: Which surfing films has Sarah Finn been the casting director for? of the 2 predicted nl questions, which is closest to the original nl question. Even if none is very close, return the one that is semantically closest? pls explain your answer as well
Appendix H Supervised Models training details
We use Hugging Face  Wolf et al. (2020), PyTorch Paszke et al. (2019) for our experiments and use the Freebase setup specified on github 555 https://github.com/dki-lab/Freebase-Setup . We use NVIDIA A100 GPU with 40 GB GPU memory and 32 GB RAM. For training the discriminator module of RetinaQA, we require 2 GPUs. (1) For the answerable experiments, we use the supervised models as specified in Patidar et al. (2024). (2) For the unanswerability experiments, we train all models from scratch. (a) We use RnG-KBQA entity linker  [^6] https://github.com/salesforce/rng-kbqa/tree/main/GrailQA/entity_linker (BSD 3-Clause License) trained on the answerable subset of GrailQAbility for all our experiments. (b) We train the RnG-KBQA path retriever on answerable subset of WebQSP [^7] https://github.com/salesforce/rng-kbqa/blob/main/WebQSP/scripts/run_ranker.sh (BSD 3-Clause License). The number of training epochs is determined by the performance of the model over the answerable questions in the dev set. (c) We train the TIARA schema retriever on the answerable subset of WebQSP  [^8] https://github.com/microsoft/KC/tree/main/papers/TIARA/src (MIT License) (d) We train the sketch generator and discriminator of RetinaQA on the answerable subset of WebQSP [^9] https://github.com/dair-iitd/RetinaQA .
Appendix I Supervised Models inference details
We use the dev set in RetinaQA, during discriminator inference for (a) determining how to best utilize the candidate paths. We might (i) not provide candidate paths (ii) provide candidate paths in GrailQA format (iii) provide candidate paths in WebQSP format. We select the best alternative based upon the performance of the model over the dev set. For the WebQSP $\rightarrow$ GrailQAbility dataset, we observe (ii) works best, whereas for the WebQSP $\rightarrow$ GraphQAbility dataset, we observe (i) works best. (b) determining the threshold value. RetinaQA applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labelled as NK. We choose the optimal value of the threshold to maximise the overall EM-s score over the dev set.

### v2

@@ 6 Conclusions
For real-world robust, low-resource KBQA, we have proposed the novel task of few-shot transfer learning with unanswerability. We have proposed FUn-FuSIC that augments for unanswerability the SoTA few-shot answerable-only KBQA transfer model by (i) iterative repair using feedback from a suite a strong and weak verifiers – including a novel back-translation based verifier – to create a set of candidate logical forms, and (ii) assessing this candidate set to detect unanswerability (and its category) and/or identify the best logical form using self consistency adapted for unanswerability.
Using two newly created datasets for this novel task, we show that FUn-FuSIC outperforms adaptations of SoTA models for this setting, and also for answerable few-shot transfer KBQA. Error analysis suggests that performing well across categories of unanswerability for few-shot transfer is still a challenge and should be a focus of further research.
@@ 8 Risks
At the highest level, our work reduces risk compared to existing KBQA systems, which when improperly adapted in a low-resource setting, incorrectly answer unanswerable questions, without admitting lack of knowledge. However, can incorrectly inferring unanswerability, citing lack of knowledge when the knowledge is in fact available, generate a new category of risk? While we cannot imagine such a risk at the present time, this may require more thought. In any case, KBQA models for unanswerability should strive to minimize this type of error, along with the other types.
@@ A.1.1 KBQA: Detailed Background
A.1.2 Few Shot Transfer Learning for KBQA
@@
A.1.3 KBQA with Unanswerability
@@
Unanswerable questions in KBQA can be categorized into (a) Schema Level Unanswerability : the question does not have a corresponding logical form that is valid for the KB , (b) Data level unanswerability: it has a valid logical form $l$ for the KB, but which on executing returns an empty answer. Schema level unanswerable questions can further be categorized into (1) Missing Class: The class/type required to construct the logical form is not defined for the KB, (2) Missing Relation: The relation required to construct the logical form is not defined for the KB, (3) Missing Topic Entity: The topic entity specified in the question is missing from the KB. Data level unanswerable questions can be categorized into (1) Missing entity: all classes and relations required to construct the logical form are present in the KB , but there exists no path from the topic entity node to the answer node in the KB due to missing intermediary entities (2) Missing Fact: all classes, relations and entities required to answer the question are present in the KB. However, the (subject, relation, object) path is not connected in the KB.
A.1.4 Few-shot Transfer Learning for KBQA with Unanswerability
@@
This setting is a generalization of few shot transfer learning for KBQA, wherein we no longer assume that the test question is answerable with respect to the Knowledge Base $G^{t}$. The training set $D^{s}$ contains thousands of answerable questions along with their corresponding logical form. The target few shots $D^{t}$ contain both answerable as well as unanswerable questions, along with their corresponding logical forms, if they exist.
A.2 Error Checks
(V2) Semantic Error (KB Inconsistency):
A syntactically correct logical form $l$ may still be inconsistent with the schema of $G$. This is a likely error even for SoTA LLMs since these are unfamiliar with the specific KB $G$. Semantic errors have different categories, such as type-incompatibility, schema hallucinations. (V2a) Incompatibility in types: $l$ contains a variable and a connecting relation whose types are incompatible in $G$. This is the case for LF1 for all three KBs in the example. (V2b) Schema hallucinations: $l$ contains schema elements (types, relations, entities) absent in $G$. (V2c) Type casting errors: Literals in $l$ are not correctly type cast for $G$, e.g. numeric literals as float for Freebase. All of these are certain checks, and are implemented using rules defined over $l$ and $G$. The feedback mentions the type of error and the specifics, e.g., the hallucinated relation, or the incompatible type-relation pair.
(V3) Question Logical form Disagreement:
FUn performs equivalence check between $l$ and $q$ using a novel multi-stage LLM pipeline. (i) The variable names in $l$ are first naturalized to $l^{n}$ considering $q$ and preserving semantics, e.g. by replacing ’?x’ with ’?actor’. (ii) $l^{n}$ is back-translated into a natural language question $q^{b}$. (iii) $q^{b}$ is finally checked for semantic equivalence with $q$. The first two steps are performed using zero-shot prompting, while the last is performed using few-shots constructed using the target few-shots $D^{t}$. More details about few-shot construction are in Sec. A.8.2. The feedback mentions lack of equivalence as the type of error.
(V4) Answer Inconsistency:
If the $l$ is syntactically and semantically correct, it is executed over $G$ to obtain an answer $a$. $a$ is then checked for compatibility with $q$. This may fail for different reasons. (V4a) $a$ (which is a set in general) contains an entity also in $l$ and therefore mentioned $q$, which is an aberration. (V4b) $a$ is empty, as in LF3 for KB2 in the example. All of these checks are implemented using rules defined over $l$ and $G$. Note that while the first two are certain checks, the last is not. An empty answer is valid for unanswerable questions, as for LF3 for KB2, but invalid for answerable ones. As before, the feedback mentions the type of error and the specifics.
A.3 EM-s: Automated check for equivalence of SPARQL programs
A.4 Performance across different categories of unanswerability
A.5 Cost Analysis
Hence, the two models are comparable in terms of cost.
A.6 FuSIC-KBQA Details
Our proposed approach FUn-FuSIC builds upon the the base architecture of FuSIC-KBQA. FuSIC-KBQA has a three step pipeline: (a) Supervised Retrieval: a supervised retriever, trained on the source domain and optionally fine-tuned on the target domain is used to obtain the top-100 classes, relations and paths that are relevant to the question asked, (c) LLM Generation: We provide the top-10 classes, top-10 relations and top-5 paths along with few-shot exemplars to generate the sparql query.
A.7 FUn-FuSIC Prompts
Here we provide details of various prompts used by FUn-FuSIC.
A.8 PUn prompt
The first prompt is for Prompting for Unanswerability (PUn).
Header Prompt Translate the following question to sparql for Freebase based on the candidate sparql, candidate entities, candidate relations and candidate entity types which are separated by "|" respectively. Please do not include any other relations, entities and entity types. Your final sparql can have three scenarios: 1. When you need to just pick from candidate sparql. 2. When you need to extend one of candidate sparql using the candidate relations and entity types. 3. When you will generate a new sparql only using the candidate entities, relations and entity types. For entity type check please use this relation "type.object.type".D o not use entity names in the query. Use specified mids. If it is impossible to construct a query using the provided candidate relations or types, return "NK". Make sure that the original question can be regenerated only using the identified entity types, specific entities and relations.
A.8.1 FUN prompt
Syntax error(V1) Feedback Correct the syntax of the following sparql query. Return ONLY the corrected sparql query without any explanation sparql: SELECT ?x AND ?y ... Virtuoso error: word AND not defined
KB Inconsistency(V2) Feedback The generated sparql has a semantic issue warning: The types of relations don’t match for variable ?x in the query. The assigned relation types by [’computer.computer_emulator.computer’, ’type.object.type computer.computer_peripheral’] are [’computer.computer’, ’computer.computer_peripheral’]. These types are mutually incompatible... Please generate again a different executable sparql using the same context and constraints. DO NOT APOLOGIZE - just return the best you can try.
Question Logical form disagreement(V3) feedback The question that you answer is NOT same as what you’ve been asked for! You have answered the question "Which opera productions has Gino Marinuzzi conducted?" but you were asked to answer "what is the name of the premiere opera production conducted by gino marinuzzi?". Please generate again a different executable sparql using the relations, classes and entities provided earlier. DO NOT APOLOGIZE - just return the best you can try.
Answer Inconsistency(V4b) feedback The generated sparql gives an empty answer when executed on freebase KG, Please generate again a different executable sparql using the same context and constraints.
Intermediate Node(V4a) feedback The generated sparql returns an intermediate type node when executed on the freebase KG. Maybe the answer node is an adjacent node to what we currently query for. Please generate again a different executable sparql using the same context and constraints.
Answer Inconsistency(V4a) feedback The logical form upon execution returns International System of Units, which is not answering the question. Please reconstruct the query using same context and constraints.
A.8.2 Prompt for Verifier V3 Question Logical Form Agreement
Naturalization of variable names(V3(i)) change the sparql query to have variable names representative of what objects they refer to. transform the variable names in this query. Do NOT change the prefix headers and relation names
@@ A.8.2 Prompt for Verifier V3 Question Logical Form Agreement
Some examples of few shots are provided below-
Question Logical Form Sgreement Check(V3(iii)) Question we answer: Who are the cricket players who have made exactly 31 stumps in one day internationals? Question originally asked: name the cricket player who has 31 odi stumps. explanation: The question we answer returns cricket players. The question originally asked also returns cricket players. The question we answer finds cricket players who have made exactly 31 stumps in one day internationals. The question originally asked also finds cricket players who have made 31 stumps in one day internationals. Both questions involve no mathematical or logical operators. Hence, they are same. Question we answer: Which game engines are successors to the Unreal Engine? Question originally asked: which video game engine’s successor engine is unreal engine? explanation: The question we answer returns game engines. The question originally asked also returns game engines. The question we answer finds successors to the Unreal Engine. The question originally asked finds the predecessor of the Unreal Engine. The reasoning steps followed by the two questions are different. Hence, they are different. [total 6 exemplars] Question we answer: Which cars drive at a speed of 80? Question originally asked: name the car with driving speed at least 80? explanation:
A.8.3 Prompt in scUn for Logical Form Selection
We store the back-translated natural language questions along with each predicted logical form. Here, we prompt the LLM to select the closest back-translated natural language question to the original question.
Best Logical Form Selection Prompt orig_nl_qn = which surf films has sarah finn served as the casting director? 1. pred_nl: Which surfing films has Sarah Finn directed the casting for? 2. pred_nl: Which surfing films has Sarah Finn been the casting director for? of the 2 predicted nl questions, which is closest to the original nl question. Even if none is very close, return the one that is semantically closest? Please explain your answer as well
A.9 Supervised Models training details
We use Hugging Face  Wolf et al. (2020), PyTorch Paszke et al. (2019) for our experiments and use the Freebase setup specified on github 777 https://github.com/dki-lab/Freebase-Setup . We use NVIDIA A100 GPU with 40 GB GPU memory and 32 GB RAM. For training the discriminator module of RetinaQA, we require 2 GPUs. (1) For the answerable experiments, we use the supervised models as specified in Patidar et al. (2024). (2) For the unanswerability experiments, we train all models from scratch. (a) We use RnG-KBQA entity linker  [^8] https://github.com/salesforce/rng-kbqa/tree/main/GrailQA/entity_linker (BSD 3-Clause License) trained on the answerable subset of GrailQAbility for all our experiments. (b) We train the RnG-KBQA path retriever on answerable subset of WebQSP [^9] https://github.com/salesforce/rng-kbqa/blob/main/WebQSP/scripts/run_ranker.sh (BSD 3-Clause License). The number of training epochs is determined by the performance of the model over the answerable questions in the dev set. (c) We train the TIARA schema retriever on the answerable subset of WebQSP  [^10] https://github.com/microsoft/KC/tree/main/papers/TIARA/src (MIT License) (d) We train the sketch generator and discriminator of RetinaQA on the answerable subset of WebQSP [^11] https://github.com/dair-iitd/RetinaQA.
A.10 Supervised Models inference details
We train all components on WebQSP, using the corresponding target domain’s dev set as a validation set for early stopping. In the absence of unanswerable questions for training, both models use a threshold fine-tuned on a dev set to detect schema-level unanswerability. We again use the target dev sets for this.
We use the dev set in RetinaQA, during discriminator inference for (a) determining how to best utilize the candidate paths. We might (i) not provide candidate paths (ii) provide candidate paths in GrailQA format (iii) provide candidate paths in WebQSP format. We select the best alternative based upon the performance of the model over the dev set. For the WebQSP $\rightarrow$ GrailQAbility dataset, we observe (ii) works best, whereas for the WebQSP $\rightarrow$ GraphQAbility dataset, we observe (i) works best. (b) determining the threshold value. RetinaQA applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK. We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
A.11 Pangu Adaptation Details
Similar to RetinaQA, We train all components on WebQSP, using the corresponding target domain’s dev set as a validation set for early stopping. We use 1 GPU for training. Same as RetinaQA, we use the dev set to determine the threshold for schema-level unanswerability. Pangu-T applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK. We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
A.12 KB-Binder and FuSIC-KBQA Adaptation Details
For KB-Binder, we make use of publicly available code  [^12] https://github.com/ltl3A87/KB-BINDER (MIT License). We use self-consistency and majority voting with 6 examples, as in the experiments in the paper. In the retrieval(-R) setting, KB-Binder samples demonstration examples by retrieving from the entire available training data. We restrict its retrieval to our target training set $D_{t}$ with 25 examples. KB-Binder reports experiments using code-davinci-002 as the LLM. For consistency and fair comparison, we replace these with gpt-4-0613 as in our approach. This model generates logical forms in s-expression, which we preserve.
For FuSIC-KBQA Patidar et al. (2024), since no code is available for this model, we use our own implementation based on the description in the paper. For FuSIC-KBQA, and FUn-FuSIC we use $temperature=0$.
A.13 Real Example of FUn
@@
A.14 Algorithm for Self Consistency with Unanswerability
The high level algorithm for self consistency with Unanswerability (scUn) is described in Algo. 3.
Algorithm 3 ScUn$(q,L,t,\mathcal{L})$
1:$(c,l,A)=\mbox{assessConf}(q,L,t,\mathcal{L})$
2:if ($\mbox{{\bf if} }(c)$, $\mbox{{\bf return}}(l,A)$
@@ A.14 Algorithm for Self Consistency with Unanswerability
The high level algorithm for assessing confidence in the set of candidate logical forms (assessConf) is described in Algo. 4.
Algorithm 4 $\mbox{assessConf}(q,L,t,\mathcal{L})$
1:$(c,L^{p},A^{p})=\mbox{popAnsNE}(L,t)$
2:if $(c)$ then
3:     $l=\mbox{selectBestNE}(q,L^{p},\mathcal{L})$
4:     $\mbox{{\bf return}}(\mbox{True},l,A^{p})$
5:end if
6:$(c,L^{p})=\mbox{popAnsE}(L,t)$
7:if $(c)$ then
8:     $l=\mbox{selectBestE}(q,L^{p},\mathcal{L})$
9:     $\mbox{{\bf return}}(\mbox{True},l,\texttt{NA})$
10:end if
[…]